# AverageTime: Enhance Long-Term Time Series Forecasting with Simple Averaging

### v3

@@ 5 Conclusion
This paper introduces AverageTime, a method that outperforms state-of-the-art Transformer architectures like TimeXer, by employing an MLP-based prediction layer and a mixed-channel embedding layer, while maintaining efficiency comparable to that of lightweight linear models. We also propose LightAverageTime, which introduces a cluster block to group highly correlated channels for joint prediction, accelerating model training. This approach significantly enhances the computational efficiency of the model while maintaining robust predictive performance. Future work could explore the underlying mechanisms of the averaging framework for long-term time series forecasting and develop more advanced approaches to enhance predictive performance.
Appendix A Hyperparameter settings
Here, we provide the detailed hyperparameter settings for AverageTime. Specifically, we list the four key hyperparameters used in our experiments on the RTX 4090D: the number of transformer layers, the number of MLP layers, the dropout rate, and whether the model is channel-independent or channel-dependent. The detailed information is shown in Table 4:
Table 4: Hyperparameters for AverageTime.
Hyperparameters Transformer layers MLP layers Dropout rate Channel id
96 0 0 0.4 1
ETTh1 192 0 1 0.6 1
336 0 2 0.8 1
720 0 2 0.8 1
96 2 1 0.85 1
ETTh2 192 3 0 0.85 1
336 2 1 0.85 1
720 2 1 0.85 1
96 0 3 0.7 1
ETTm1 192 0 3 0.3 1
336 0 3 0.3 1
720 0 3 0.3 1
96 0 3 0.5 1
ETTm2 192 0 3 0.5 1
336 0 3 0.5 1
720 0 3 0.5 1
96 0 1 0 1
Weather 192 0 1 0 1
336 0 1 0 1
720 0 1 0.7 1
96 1 3 0 1
Solar 192 1 3 0 1
336 1 3 0 1
720 1 3 0 1
96 2 3 0 1
ECL 192 2 3 0 1
336 2 3 0 1
720 3 3 0 1
96 3 0 0 0
Traffic 192 3 0 0 0
336 3 0 0 0
720 3 0 0 0
Since hyperparameters play a crucial role in model performance, it is important to make them accessible and provide guidance on how to adjust them effectively.
For the ETTh2, Solar, and ECL datasets, we use both transformer encoders and MLPs for channel embedding. For ETTh1, ETTm1, ETTm2, and Weather, we use only MLP layers for channel embedding. For the Traffic dataset, we use only transformer encoders and adopt a channel-dependent approach to improve efficiency (though this is not necessary).
For datasets with fewer channels, such as the ETT series and Weather datasets, MLPs are sufficient for channel embedding. In contrast, for datasets with a larger number of channels, transformer encoders are preferable for enhancing model performance. Additionally, as the number of channels increases, a lower dropout rate should be set for better model stability.
Appendix B Detail information for long time series forecasting result
The detailed information from Table 2 is presented in Table 5:
Table 5: Detail result for AverageTime and LightAverageTime in Long-term time series forecasting.
Datasets Avg.Time LAvg.Time TimeXer CycleNet iTrans. RLinear PatchTST TimesNet
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.378 0.393 0.376 0.392 0.382 0.403 0.375 0.395 0.386 0.405 0.386 0.395 0.414 0.419 0.384 0.402
ETTh1 192 0.434 0.424 0.435 0.425 0.429 0.435 0.436 0.428 0.441 0.436 0.437 0.424 0.460 0.445 0.436 0.429
336 0.475 0.446 0.474 0.445 0.468 0.448 0.496 0.455 0.487 0.458 0.479 0.446 0.501 0.466 0.491 0.469
720 0.466 0.461 0.464 0.460 0.469 0.461 0.520 0.484 0.503 0.491 0.481 0.470 0.500 0.488 0.521 0.500
96 0.290 0.340 0.292 0.343 0.286 0.338 0.298 0.344 0.297 0.349 0.288 0.338 0.302 0.348 0.340 0.374
ETTh2 192 0.370 0.390 0.376 0.393 0.363 0.389 0.372 0.396 0.380 0.400 0.374 0.390 0.388 0.400 0.402 0.414
336 0.414 0.426 0.417 0.428 0.414 0.423 0.431 0.439 0.428 0.432 0.415 0.426 0.426 0.433 0.452 0.452
720 0.413 0.435 0.423 0.441 0.408 0.432 0.450 0.458 0.427 0.445 0.420 0.440 0.431 0.446 0.462 0.468
96 0.314 0.353 0.318 0.358 0.318 0.356 0.319 0.360 0.334 0.368 0.355 0.376 0.329 0.367 0.338 0.375
ETTm1 192 0.356 0.375 0.360 0.378 0.362 0.383 0.360 0.381 0.387 0.391 0.391 0.392 0.367 0.385 0.374 0.387
336 0.392 0.398 0.392 0.400 0.395 0.407 0.389 0.403 0.426 0.420 0.424 0.415 0.399 0.410 0.410 0.411
720 0.463 0.438 0.460 0.435 0.452 0.441 0.447 0.441 0.491 0.459 0.487 0.450 0.454 0.439 0.478 0.450
96 0.170 0.251 0.170 0.253 0.171 0.256 0.163 0.246 0.180 0.264 0.182 0.265 0.175 0.259 0.187 0.267
ETTm2 192 0.235 0.294 0.236 0.295 0.237 0.299 0.229 0.290 0.250 0.309 0.246 0.304 0.241 0.302 0.249 0.309
336 0.293 0.332 0.297 0.334 0.296 0.338 0.284 0.327 0.311 0.348 0.307 0.342 0.305 0.343 0.321 0.351
720 0.396 0.393 0.396 0.393 0.392 0.394 0.389 0.391 0.412 0.407 0.407 0.398 0.402 0.440 0.408 0.403
96 0.156 0.203 0.156 0.203 0.157 0.205 0.158 0.203 0.174 0.214 0.192 0.232 0.177 0.218 0.172 0.220
Weather 192 0.205 0.248 0.205 0.245 0.204 0.247 0.207 0.247 0.221 0.254 0.240 0.271 0.225 0.259 0.219 0.261
336 0.263 0.289 0.260 0.286 0.261 0.290 0.262 0.289 0.278 0.296 0.292 0.307 0.278 0.297 0.280 0.306
720 0.345 0.345 0.343 0.342 0.340 0.341 0.344 0.344 0.358 0.349 0.364 0.353 0.354 0.348 0.365 0.359
96 0.198 0.247 0.215 0.257 0.201 0.259 0.186 0.245 0.203 0.237 0.322 0.339 0.225 0.270 0.250 0.292
Solar 192 0.227 0.270 0.234 0.270 0.229 0.276 0.201 0.256 0.233 0.261 0.360 0.358 0.253 0.288 0.296 0.318
336 0.241 0.281 0.243 0.283 0.251 0.293 0.228 0.276 0.248 0.273 0.397 0.369 0.270 0.298 0.319 0.330
720 0.251 0.286 0.255 0.281 0.251 0.295 0.226 0.274 0.249 0.275 0.396 0.361 0.269 0.299 0.338 0.337
96 0.142 0.240 0.144 0.242 0.140 0.242 0.136 0.229 0.148 0.240 0.201 0.281 0.195 0.285 0.168 0.272
ECL 192 0.161 0.256 0.163 0.258 0.157 0.256 0.152 0.244 0.162 0.253 0.201 0.283 0.199 0.289 0.184 0.289
336 0.177 0.275 0.181 0.278 0.176 0.275 0.170 0.264 0.178 0.269 0.215 0.298 0.215 0.305 0.198 0.300
720 0.217 0.310 0.217 0.310 0.211 0.306 0.212 0.299 0.225 0.317 0.257 0.331 0.256 0.337 0.220 0.320
96 0.394 0.263 0.396 0.265 0.428 0.271 0.458 0.296 0.395 0.268 0.649 0.389 0.462 0.295 0.593 0.321
Traffic 192 0.415 0.274 0.418 0.276 0.448 0.282 0.457 0.294 0.417 0.276 0.601 0.366 0.466 0.296 0.617 0.336
336 0.428 0.280 0.430 0.281 0.473 0.289 0.470 0.299 0.433 0.283 0.609 0.369 0.482 0.304 0.629 0.336
720 0.457 0.297 0.460 0.300 0.516 0.307 0.502 0.314 0.467 0.302 0.647 0.387 0.514 0.322 0.640 0.350
Our model shows great robotness for all datasets, showing the effectivness of average framework for long-term time series forecasting.

### v4

@@ 5 Conclusion
This paper presents AverageTime, which surpasses the state-of-the-art Transformer-based architecture TimeXer in prediction accuracy by extracting and fusing channel-wise information, while maintaining a lightweight model structure. AverageTime exhibits considerable flexibility in its architecture, with its core design aiming to acquire and integrate richer information. To validate this idea, we further introduce sequence decomposition information into the model, achieving consistent performance gains without additional hyperparameter tuning. We also propose LightAverageTime, which incorporates a community clustering algorithm to allow highly similar channels to share parameters. This approach significantly accelerates both training and inference speeds while largely preserving model performance. Future work could explore providing more effective information for time series fusion and designing efficient fusion mechanisms to further enhance the model’s predictive capability.